# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v5

@@ 1 Introduction
Going beyond single orbital models, it is natural to consider multi-orbital extensions which are relevant for real materials. In such multi-orbital models, the complexity of the orbital structure introduces new competing energy scales, as exemplified by the Hund’s metals [52, 53, 54, 55, 56] which turn out to be relevant for a large variety of materials including iron-based superconductors [57, 58, 59, 54, 60], ruthenates [61, 62] and molybdates [63], to mention a few. This more intricate local structure of the interaction tensor has profound consequences for a number of numerical approaches, e.g. leading to a sign problem in some QMC variants [64] and, in general, to a larger numerical cost. Thus, extending numerical techniques to multi-orbital systems poses often not only an analytical but also a computational challenge which has to be overcome. Furthermore, approximations known to work well in the single orbital case are not guaranteed to be equally adequate for multi-orbital systems.
@@
One very successful numerical method for systems with weak to intermediate correlations is the two particle self-consistent (TPSC) approach [80, 81, 82], which was first formulated for the single-band Hubbard model [83, 84], and subsequently extended to multi-site [85, 86], non-$SU(2)$ [87, 88], multi-orbital [89, 66, 90, 91] and non-equilibrium [92, 93] problems. Furthermore, it has been combined with DMFT to extend its range of validity [94, 95, 96]. TPSC has been applied extensively to single orbital models [97, 98, 80, 99, 100, 101, 102, 103, 104], for which it yields remarkably accurate results [50] at a comparably low numerical cost. Its combination with DMFT does not require the calculation of a vertex, but only the two-particle density matrix, which can be evaluated at much lower numerical cost. If this methodology works reliably in multi-orbital systems, TPSC and TPSC+DMFT will become prime contenders for the study of complex correlated materials.
@@ 2 Multi-orbital TPSC
In this chapter we present two variants of the multi-orbital TPSC formalism. The derivation follows closely analogous derivations for the single orbital case [80, 82]. In the following we will consider a general tight-binding Hamiltonian with local and instantaneous interactions
@@ 3.1 Model with pure density-density interactions
$\chi^{0}_{o_{1},o_{2},o_{3},o_{4}}(\boldsymbol{q},\omega)=G_{o_{1}o_{4}}(\boldsymbol{k},\nu)G_{o_{2}o_{3}}(\boldsymbol{k}-\boldsymbol{q},\nu-\omega)=\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\chi^{0}(\boldsymbol{q},\omega)\,,$ (51)
@@
${H}_{\text{int}}=\sum_{i,o_{1},o_{2},s}Un_{i,o_{1},s}n_{i,o_{2},\bar{s}}+(1-\delta_{o_{1}o_{2}})Un_{i,o_{1},s}n_{i,o_{2},{s}}\;.$ (52)
This interaction in combination with the kinetic term leads to a vanishing $\Sigma^{0}$ (see App. (A.2) Eq. 99 or Ref. [91]). Since $J=0$ there is no spin selectivity and we thus expect $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}$. Furthermore, since $U=U^{\prime}$, there is no energy difference for double occupying the same or different orbitals, and hence $\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$. This also implies that the double occupancy $\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$ should decrease less rapidly with $U$ than in the case with $U^{\prime}=0$ (which then is equivalent to the single orbital Hubbard model). With this in mind we now turn to TPSC.
@@
We note that the interacting susceptibility is identical to the one of a single orbital model with the same renormalized interaction value. The only non-zero components of the susceptibility are $\chi_{1111}$ and $\chi_{1221}$. Thus, from Eq. (39) and Eq. (40) it immediately follows that $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$, as we expected from the arguments above. Inserting these findings into the Ansatz equations yields
@@
Due to the symmetries we found, the right hand side is orbitally independent and gives the same equation for the renormalized interactions as the single orbital case [80, 82]. This result is also independent of what filling we initially chose and is a general feature of the considered model.
Therefore, in contrast to our expectations, the specific form of the susceptibility implies that the converged TPSC loop will result in the same double occupancy as the one for the one-orbital model. In other words, adding an inter-orbital interaction of the same strength as the bare interaction does not make a difference for the double occupation predicted by the method. This is in contrast to the physically expected result discussed above – in the single orbital model in the strong coupling limit, we would expect zero double occupancy. However in two-orbital model, different states in which no orbital is doubly occupied and states in which one orbitals is doubly occupied are degenerate in energy, so the double occupancy should decay more slowly.
@@
Figure 2: (a) Comparison between TPSC and ED at $\beta=2$ for different $U$ at $J=0$. We observe the same phenomenology as expected from our analytical analysis. (b) Time dependence of the density-density correlator for a single (SO) and a two-orbital (MO) two-site Hubbard model. No enhancement of the time dependence is visible, and hence this cannot explain the larger deviation to DMFT in the multi-orbital case.
Let us first answer the second question. In the single-orbital case we fix the local and static expectation values to be the exact ones as determined by the sum rules. Crucially, the single site contains no further internal structure meaning that the exact local expectation value is always proportional to the non-interacting one. Thus representing this expectation value locally by a renormalized Hartree-Fock expectation value works, i.e., we can always get the correct value of the double occupancy by multiplying the non-interacting double occupancy with a single number. Multi-orbital systems have additional degrees of freedom, which enter the local and static contribution. We still approximate the TPEVs in Eq. (8) by a Hartree-Fock decoupling. In other words, we approximate the ground state by a single Slater-determinant. However, the local Hilbert-space is now spanned by more than a single wave-function and all of these states contribute to the expectation value. Thus, introducing interactions can fundamentally alter the structure of the TPEV’s such that these can deviate from a behavior multiplicably representable by the single Slater determinant from Hartree-Fock, which explains the issue we described in the first question posed above. As an example for this, let us consider a single dimer (i.e. a single site in Fig. 1) at half-filling with a Hubbard-Kanamori interaction. For simplicity, let us call the two orbitals $a$ and $b$. The Hamiltonian then contains only interaction terms and reads
@@
This we can readily diagonalize in the subspace containing two electrons (|$\ket{\uparrow_{a},\downarrow_{a}}$, |$\ket{\uparrow_{a},\downarrow_{b}}$, |$\ket{\uparrow_{b},\downarrow_{a}}$, |$\ket{\uparrow_{a},\uparrow_{b}}$, |$\ket{\downarrow_{a},\downarrow_{b}}$, |$\ket{\uparrow_{b},\downarrow_{b}}$). We find three different groups of eigenstates: The first one is three-fold degenerate with eigenvalue $U-3J$ and spanned by |$\ket{\uparrow_{a},\uparrow_{b}}$, |$\ket{\downarrow_{a},\downarrow_{b}}$ and (|$(\ket{\uparrow_{a},\downarrow_{b}}-\ket{\uparrow_{b},\downarrow_{a}})/\sqrt{2}$. These three states form the $S=1$ sector. Next, we do have two states with eigenvalue $U-J$, (|$(\ket{\uparrow_{a},\downarrow_{b}}+\ket{\uparrow_{b},\downarrow_{a}})/\sqrt{2}$ and (|$(\ket{\uparrow_{a},\downarrow_{a}}-\ket{\uparrow_{b},\downarrow_{b}})/\sqrt{2}$. And lastly, a single eigenstate with eigenvalue $U+J$, (|$(\ket{\uparrow_{a},\downarrow_{a}}+\ket{\uparrow_{b},\downarrow_{b}})/\sqrt{2}$.
Therefore, in the absence of $J$ all six eigenvalues are degenerate in the local problem. In Hartree-Fock (and by inheritance in TPSC) only one of these six states is represented by the decoupling and the others are missed entirely (thus behaving like the single orbital model in the absence of $J$). Thereby, the basic idea of TPSC, which relies on fixing the local and static expectation values starting from Hartree-Fock, breaks down due to the inadequateness of Hartree-Fock. Furthermore, since this issue seems to be inherited from the Hartree-Fock decoupling, which is central to TPSC, there is no simple repair for this problem without fundamentally altering the structure of the method.
@@ 3.2 Analysis of the model with $J=U/3$
Since we found in the last section that the problems arising can be traced back to the Hartree-Fock ansatz, the question arises whether we can again understand the improvement as an inherited effect. To test this, we again consider the half-filled dimer described above. The three groups of eigenvalues we found above split with increasing $J$. Thus, while the point $U=3J$ is not special, the separation between the excited state and the ground state sectors grows linearly with $J$. Furthermore, the $S=1$ subspace is special, as two states only contribute to a single TPEV ($\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$) and the other state essentially behaves like the single orbital model. Therefore, one should be able to approximate the TPEV’s by a scaling in all components but the $\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$ one, which is incorrectly estimated because of the missing degenerate states. This is exactly what is observed for both TPSC5 and in Ref. [90]. Thus we conclude that in the end this improvement is related to the structure and number of states determining the local and static expectation values. Furthermore, the improved performance is entirely explained by the improved performance of the Hartree-Fock ansatz.
@@
From these observations, we can extract a guiding principle of when to expect TPSC to work reliably in a multi-orbital setting and when not: If there is a gap in the local spectrum between the ground state and the excited states, TPSC should perform better than if the states are close in energy. Furthermore, when analyzing the nature of the states, we can extract which TPEV is expected to deviate strongly and which not. This is intrinsically tied to Hartree-Fock being essentially a ground state targeting approach which performs better at lower temperatures. However, at very low temperatures, we expect the envelope diagram to give a sizable contribution near the Fermi-level inducing a non-local and non-static irreducible Vertex [107, 108]. Additionally, it was shown in an RG analysis that in the low temperature limit, the vertex is only RG relevant at the Fermi-surface again implying a momentum and Frequency dependence strongly confined in both spaces [28]. Therefore, the local and static approximation of the vertex becomes more and more inappropriate at lower temperatures which is why TPSC is expected to fail at very low temperatures.
@@
In summary, the regimes in which multi-orbital TPSC is guaranteed to work well are the weak coupling limit and a small region in which excited states do not contribute to the local TPEVs (controlled by temperature and the gap), while the vertex still is approximately local and static. The former criterion is accessible by inspection, while the latter one is not clear a priori and requires other numerical simulations to be fully gauged. Notably, the issue described above is partially resolved when starting from DMFT TPEVs [91, 95], since fixing the double occupancies through an external source prevents the negative feedback from a wrong starting point.
@@ 3.3.2 Comparison to D-TRILEX - self-energy and susceptibilities
A similar picture emerges for the susceptibility: all TPSC variants over(under)-estimate the spin(charge) susceptibility in comparison to D-TRILEX, in agreement with what was reported in Ref. [95]. However, the overestimation is weaker for TPSC5 and stronger for TPSC3. Most notably, the charge susceptibility is even more strongly suppressed than in prior TPSC formulations. This disagreement between the TPSC variants is a consequence of the vertex having a different from, see App. B.1, which leads to problems in the absence of $J$, see Ref. [90] but better agreement to the reference data at large $J$, see Fig. 3.
@@ 4 Conclusions and Outlook
In this paper we introduced two variants of a fully self-consistent multi-orbital TPSC approach dubbed TPSC3 and TPSC5. We analyzed the structure of the equations and showed analytically that in the limit of vanishing Hund’s coupling the Ansatz equations result in the same TPEVs as for the single orbital model, which is unphysical. Furthermore, we provided an explanation for this behavior in terms of the equivalence between the Hatree-Fock decomposition and a minimization in terms of single Slater determinants, highlighting that the influences from many low-energy states onto the local TPEVs cannot be fully reproduced by a single state with modified prefactors. By considering the local Hilbert space we also provided an understanding of why the approach performs better at larger $J/U$. Notably, this understanding allows to assess the expected quality of the results from TPSC by inspecting the local spectrum. On a qualitative level, we found that the TPEVs from TPSC5 outperform the ones from TPSC3 when compared to DMFT. In the limiting case of $J=U/3$ the new variants do not outperform the conceptually simpler but approximate approach suggested in Ref. [90].

### v7

@@ 1 Introduction
Going beyond single-orbital models, it is natural to consider multi-orbital extensions which are relevant for real materials. In such multi-orbital models, the complexity of the orbital structure introduces new competing energy scales, as exemplified by the Hund’s metals [52, 53, 54, 55, 56] which turn out to be relevant for a large variety of materials including iron-based superconductors [57, 58, 59, 54, 60], ruthenates [61, 62] and molybdates [63], to mention a few. This more intricate local structure of the interaction tensor has profound consequences for a number of numerical approaches, e.g. leading to a sign problem in some QMC variants [64] and, in general, to a larger numerical cost. Thus, extending numerical techniques to multi-orbital systems poses often not only an analytical but also a computational challenge which has to be overcome. Furthermore, approximations known to work well in the single-orbital case are not guaranteed to be equally adequate for multi-orbital systems.
@@
One very successful numerical method for systems with weak to intermediate correlations is the two particle self-consistent (TPSC) approach [80, 81, 82], which was first formulated for the single-band Hubbard model [83, 84], and subsequently extended to multi-site [85, 86], non-$SU(2)$ [87, 88], multi-orbital [89, 66, 90, 91] and non-equilibrium [92, 93] problems. Furthermore, it has been combined with DMFT to extend its range of validity [94, 95, 96]. TPSC has been applied extensively to single-orbital models [97, 98, 80, 99, 100, 101, 102, 103, 104], for which it yields remarkably accurate results [50] at a comparably low numerical cost. Its combination with DMFT does not require the calculation of a vertex, but only the two-particle density matrix, which can be evaluated at much lower numerical cost. If this methodology works reliably in multi-orbital systems, TPSC and TPSC+DMFT will become prime contenders for the study of complex correlated materials.
@@ 2 Multi-orbital TPSC
In this chapter we present two variants of the multi-orbital TPSC formalism. The derivation follows closely analogous derivations for the single-orbital case [80, 82]. In the following we will consider a general tight-binding Hamiltonian with local and instantaneous interactions
@@ 3.1 Model with pure density-density interactions
$\chi^{0}_{o_{1},o_{2},o_{3},o_{4}}(\boldsymbol{q},\omega)=\sum_{\boldsymbol{k},\nu}G_{o_{1}o_{4}}(\boldsymbol{k},\nu)G_{o_{2}o_{3}}(\boldsymbol{k}-\boldsymbol{q},\nu-\omega)=\delta_{o_{1},o_{4}}\delta_{o_{2},o_{3}}\chi^{0}(\boldsymbol{q},\omega)\,,$ (51)
@@
${H}_{\text{int}}=\sum_{i,o_{1},o_{2},s}Un_{i,o_{1},s}n_{i,o_{2},\bar{s}}+(1-\delta_{o_{1}o_{2}})U^{\prime}n_{i,o_{1},s}n_{i,o_{2},{s}}\;.$ (52)
This interaction in combination with the kinetic term leads to a vanishing $\Sigma^{0}$ (see App. (A.2) Eq. 99 or Ref. [91]). Since $J=0$ there is no spin selectivity and we thus expect $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}$. Furthermore, since $U=U^{\prime}$, there is no energy difference for double occupying the same or different orbitals, and hence $\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$. This also implies that the double occupancy $\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$ should decrease less rapidly with $U$ than in the case with $U^{\prime}=0$ (which then is equivalent to the single-orbital Hubbard model). With this in mind we now turn to TPSC.
@@
We note that the interacting susceptibility is identical to the one of a single-orbital model with the same renormalized interaction value. The only non-zero components of the susceptibility are $\chi_{1111}$ and $\chi_{1221}$. Thus, from Eq. (39) and Eq. (40) it immediately follows that $\braket{n_{0}^{\uparrow}n_{1}^{\uparrow}}=\braket{n_{0}^{\uparrow}n_{1}^{\downarrow}}=\braket{n_{0}^{\uparrow}n_{0}^{\downarrow}}$, as we expected from the arguments above. Inserting these findings into the Ansatz equations yields
@@
Due to the symmetries we found, the right hand side is orbitally independent and gives the same equation for the renormalized interactions as the single-orbital case [80, 82]. This result is also independent of what filling we initially chose and is a general feature of the considered model.
Therefore, in contrast to our expectations, the specific form of the susceptibility implies that the converged TPSC loop will result in the same double occupancy as the one for the one-orbital model. In other words, adding an inter-orbital interaction of the same strength as the bare interaction does not make a difference for the double occupation predicted by the method. This is in contrast to the physically expected result discussed above – in the single-orbital model in the strong coupling limit, we would expect zero double occupancy. However in the two-orbital model, different states in which no orbital is doubly occupied and states in which one orbitals is doubly occupied are degenerate in energy, so the double occupancy should converge towards a nonzero value.
@@
Figure 2: (a) Comparison between TPSC and ED at $\beta=2$ for different $U$ at $J=0$. We observe the same phenomenology as expected from our analytical analysis. (b) Time dependence of the density-density correlator for a single (SO) and a two-orbital (MO) two-site Hubbard model from ED at $\beta=2$. No enhancement of the time dependence is visible, and hence this cannot explain the larger deviation to DMFT in the multi-orbital case.
Let us first answer the second question. In the single-orbital case we fix the local and static expectation values to be the exact ones as determined by the sum rules. Crucially, the single site contains no further internal structure meaning that the exact local expectation value is always proportional to the non-interacting one. Thus representing this expectation value locally by a renormalized Hartree-Fock expectation value works, i.e., we can always get the correct value of the double occupancy by multiplying the non-interacting double occupancy with a single number.
Multi-orbital systems have additional degrees of freedom, which enter the local and static contribution. However, we still approximate the TPEVs in Eq. (8) by a Hartree-Fock decoupling. To make matters simple, let us first discuss the zero temperature limit. In this limit, the perturbative many-body theory Hartree-Fock variant we use to decouple the expectation values becomes equivalent to the variational Hartree-Fock variant. Thus, let us consider the ladder one for now. Variational Hartree-Fock approximates the ground-state by a single-Slater determinant. However, in cases in which the ground state is degenerate, all of the degenerate states of the local Hilbert-space influence the expectation values. Thereby, the TPEV’s we obtain from Hartree-Fock are biased by the introduction of an artificial preference to a single state which was selected arbitrarily. Of course, TPSC is a finite temperature method - so while this argument explains what we observe at sufficiently low temperatures, it can only give us a hint about the physical origin. To better understand the issue let us consider a single dimer (i.e. a single site in Fig. 1) at half-filling with a Hubbard-Kanamori interaction. For simplicity, let us call the two orbitals $a$ and $b$. The Hamiltonian then contains only interaction terms and reads
@@
This Hamiltonian can be readily diagonalized in the subspace containing two electrons (|$\ket{\uparrow_{a},\downarrow_{a}}$, |$\ket{\uparrow_{a},\downarrow_{b}}$, |$\ket{\uparrow_{b},\downarrow_{a}}$, |$\ket{\uparrow_{a},\uparrow_{b}}$, |$\ket{\downarrow_{a},\downarrow_{b}}$, |$\ket{\uparrow_{b},\downarrow_{b}}$). We find three different groups of eigenstates: The first one is three-fold degenerate with eigenvalue $U-3J$ and spanned by |$\ket{\uparrow_{a},\uparrow_{b}}$, |$\ket{\downarrow_{a},\downarrow_{b}}$ and (|$(\ket{\uparrow_{a},\downarrow_{b}}-\ket{\uparrow_{b},\downarrow_{a}})/\sqrt{2}$. These three states form the $S=1$ sector. Next, we do have two states with eigenvalue $U-J$, (|$(\ket{\uparrow_{a},\downarrow_{b}}+\ket{\uparrow_{b},\downarrow_{a}})/\sqrt{2}$ and (|$(\ket{\uparrow_{a},\downarrow_{a}}-\ket{\uparrow_{b},\downarrow_{b}})/\sqrt{2}$. And lastly, a single eigenstate with eigenvalue $U+J$, (|$(\ket{\uparrow_{a},\downarrow_{a}}+\ket{\uparrow_{b},\downarrow_{b}})/\sqrt{2}$.
In this specific case, in the absence of $J$ we directly observe that the local self-energy, see Eq. 17, becomes a diagonal matrix independent of the orbital, as the Fock term vanishes (this is also true for the case discussed in Sec. 3). Thus, fixing the particle number absorbs the Hartree-Fock correction into the chemical potential. Therefore, the Hartree-Fock solution at fixed particle number is the same as the non-interacting model - directly explaining why above we observed that adding the inter-orbital interactions was reproducing the single-orbital case. Since the TPSC ansatz in this limit starts from a point indistinguishable from the single-orbital case due to the approximations made, it generates the same correction and follows the same trajectory.
@@ 3.2 Analysis of the model with $J=U/3$
Since we found in the last section that the problems arising can be traced back to the Hartree-Fock ansatz, the question is whether we can again understand the improvement as an inherited effect. To test this, we again consider the half-filled dimer described above. At $T=0$ we again have to consider degeneracies: We found three groups of eigenvalues above, whose eigenvalues split with increasing $J$. Thus, while the point $U=3J$ is not special, the separation between the excited states and the ground state sector grows linearly with $J$. Furthermore, the $S=1$ subspace is special, as two states only contribute to a single TPEV ($\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$) and the other state essentially behaves like the single-orbital model. Thus, at $T=0$ we would expect already better performance of the ansatz as the degeneracy is lifted. Therefore, one should be able to approximate the exact TPEV’s by a scaling in all components but the $\braket{n^{\uparrow}_{0}n^{\uparrow}_{1}}$ one, which is incorrectly estimated because of the degenerate states being incorrectly represented in the perturbation theory. This is exactly what is observed for both TPSC5 and in Ref. [90]. How is this lifting of the ground state degeneracy influencing the behavior at nonzero temperature? - At the first glance, it does not directly, as still the introduction of the Hund’s coupling merely changes the numerical value of the self-energy but it still remains a diagonal matrix without any orbital selectivity. Therefore, the starting point is again the non-interacting model, just as before. However, introducing $J$ allows for a differentiation of equal orbital and unequal orbital cases leading to a better correction within TPSC. Thus we conclude that in the end the improvement is related to the structure of the equation changing by the introduction of a Hund’s coupling. Further, we note that we can analogously understand the improvement by considering the number of degenerate ground-states within the local Hilbert-space  [^1] That this analogy works so well is a consequence of the true many-body Green’s function being better representable by Hartree-Fock in the limit of a gapped ground state..
@@
From these observations and the connection at $T=0$ to variational Hartree-Fock, we can extract guiding principles of when to expect TPSC to work reliably in a multi-orbital setting and when not: If there is a gap in the local spectrum between the ground state and the excited states, TPSC should perform better than if the states are closer in energy. Furthermore, when analyzing the nature of the states, we can extract which TPEV is expected to deviate strongly and which not. In other words, in cases in which Hartree-Fock produces a sensible starting point, TPSC is expected to perform well. Whether this is the case or not strongly depends on the validity of approximating the self-energy to be static. On the other hand, at very low temperatures, we expect the envelope diagram to give a sizable contribution near the Fermi-level inducing a non-local and non-static irreducible vertex [107, 108]. Additionally, it was shown in an RG analysis that in the low temperature limit, the vertex is only RG relevant at the Fermi-surface again implying a momentum and Frequency dependence strongly confined in both spaces [28]. Therefore, the local and static approximation of the vertex becomes more and more inappropriate at lower temperatures which is why TPSC is expected to fail at very low temperatures.
@@
In summary, the regimes in which multi-orbital TPSC is guaranteed to work well are the region in which Hartree-Fock in itself provides a sensible starting point, e.g. the weak-coupling limit and cases in which the self-energy is well approximated by a static function. Furthermore, the vertex has to be approximately local and static. The former criterion we can try to gauge by considering the spectrum of the local Hilbert-space due to the connection to the variational Hartree-Fock flavor. The latter one is not clear a priori and requires other numerical simulations to be fully gauged. Notably, the issue described above is partially resolved when starting from DMFT TPEVs [91, 95], since fixing the double occupancies through an external source prevents the negative feedback from a wrong starting point.
@@ 3.3.2 Comparison to D-TRILEX - self-energy and susceptibilities
A similar picture emerges for the susceptibility, see Fig. 6: all TPSC variants over(under)-estimate the spin(charge) susceptibility in comparison to D-TRILEX, in agreement with what was reported in Ref. [95]. However, the overestimation is weaker for TPSC5 and stronger for TPSC3. Most notably, the charge susceptibility is even more strongly suppressed than in prior TPSC formulations. This disagreement between the TPSC variants is a consequence of the vertex having a different from, see App. B.1, which leads to problems in the absence of $J$, see Ref. [90] but better agreement to the reference data at large $J$, see Fig. 3.
@@ 4 Conclusions and Outlook
In this paper we introduced two variants of a fully self-consistent multi-orbital TPSC approach dubbed TPSC3 and TPSC5. We analyzed the structure of the equations and showed analytically that in the limit of vanishing Hund’s coupling the Ansatz equations result in the same TPEVs as for the single-orbital model, which is unphysical. Furthermore, we provided an explanation for this behavior in terms of the underlying Hartree-Fock decomposition and its connection to the variational Hartree-Fock approach, as well as the general criteria under which a static self-energy is expected to be sufficiently accurate. By considering the local Hilbert space we also provided an understanding of why the approach performs better at larger $J/U$. Notably, this understanding allows to assess the expected quality of the results from TPSC by inspecting the local spectrum. On a qualitative level, we found that the TPEVs from TPSC5 outperform the ones from TPSC3 when compared to DMFT. In the limiting case of $J=U/3$ the new variants do not outperform the conceptually simpler but approximate approach suggested in Ref. [90].